# Q-Net: Transferable Queue Length Estimation via Kalman-based Neural Networks

### v1

@@ 6.2 Performance
Figure 7: Box plots of absolute estimation errors for baseline models (OSD, ISC) and Q-Net across all test days for section N1-IN. All p-values between the error distributions of baseline methods and our approach are less than $10^{-5}$, indicating statistically significant differences.
Rotterdam is the second-largest city in the Netherlands. To evaluate our method, we first filter out road segments lacking sufficient data availability, either due to contractual restrictions with the municipality or the presence of significant missing values. Among the remaining candidates, we select N1-IN as it is one of the major urban road sections in the Rotterdam traffic network. Evaluating our model on this section is of practical importance because it directly connects to the tunnel in Rotterdam, a critical point of interest for the municipality’s traffic management efforts. A satellite view of the section is shown in Figure 6. After vehicles exit the upstream intersection, they can either enter the ring interchange or continue through N1-IN, leading to significant unobserved traffic flow exiting the study section.
Table 3: Performance comparison between the proposed model and baseline models in terms of RMSE, MAE, and MAPE metrics.
@@
OSD 214.62 130.57 104.25 18.74 50.15 130.32 247.09 308.21 69.96
ISC 188.99 127.29 95.78 34.20 68.41 130.38 245.97 269.03 59.74
Q-Net (ours) 71.65 49.51 48.60 5.29 36.94 79.43 26.79 68.04 17.91
Improvement 62.09% 61.10% 49.26% 71.77% 26.34% 39.05% 89.11% 74.71% 70.02%
In Table 3, we evaluate model performance on the test days using three metrics: Root Mean Square Error (RMSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE). For MAPE, we consider only instances where the ground truth queue length exceeds 10 meters to avoid skewed results from disproportionately large percentage errors on near-zero values. Considering that traffic dynamics vary substantially across time periods, metrics are reported for the entire day, the morning peak, and the afternoon peak. The morning and afternoon peak hours are defined according to the guidelines from the Dutch Ministry of Infrastructure and Water Management (Mobiliteitsscan, ).
Our results show that the proposed Q-Net consistently outperforms the baseline approaches across all metrics and time periods. We observe that MAPE is generally higher during the morning peak hours. Upon reviewing the ground truth data, we found that substantial queues during the morning peak occurred only on 2023-11-14, whereas the other test days exhibited much shorter queues. As MAPE is sensitive to low ground truth values, shorter queues naturally result in higher percentage errors.
@@
Figure 8: The aFCD visualization and queue length estimation comparison among baselines (OSD and ISC), and our proposed method for section N1-IN on 2023-11-14.
Figure 7 shows the box plot of absolute estimation errors for each method across all test days. Q-Net demonstrates significantly lower mean errors and reduced variance. Notably, all methods show smaller errors on 2023-11-18, which coincides with a Saturday. This improvement is likely due to lower traffic demand on weekends, resulting in shorter queues that are easier to predict.
We visualized the estimation results for 2023-11-14 in Figure 8. The results indicate that both baselines are highly sensitive to errors in the aFCD data. As shown in Figure 8(a), these errors manifest in several ways: (a) during the congested period (15:00–19:00), relatively high speeds are recorded within the queue; (b) the first FCD segment (closest to the stop line) often reports higher speeds despite the presence of long queues, likely due to sparse probe vehicle coverage or smoothing artifacts introduced during aggregation; and (c) isolated low-speed patches appear in the middle of the segment, which do not align with the actual queue positions observed in the ground truth.
These inconsistencies in the aFCD signal lead to degraded performance in the baseline methods. Baseline OSD exhibits sharp, erratic fluctuations and fails to capture the afternoon peak queue. Baseline ISC, which relies on iso-speed contours, suffers from discontinuities and generates multiple queue estimates for a single time step, highlighting its sensitivity to noisy input data. In contrast, Q-Net successfully captures both the trend and variations in queue length, even under noisy input conditions, thanks to KalmanNet’s ability to dynamically track noise covariance. Furthermore, when we focus on the period from 13:30 to 14:00, we observe a queue formation delay in the aFCD. Despite this, Q-Net accurately tracks the onset of the queue, likely due to the initial estimate derived from the loop detector cumulative count data, which helps mitigate the lag present in the FCD signal.
@@ 6.3 Transferability
This subsection focuses on the transferability analysis of our method. We introduce two additional sections from the Rotterdam urban area that are adjacent to N1-IN: N3-IN and N3-OUT, as shown in Figure 9. Both sections are located underground and feature an above-ground split after vehicles exit their upstream intersections. These diverging flows are eventually merged at the downstream intersection. The upstream intersection of N3-IN includes lanes with combined directions (through and right-turn), meaning unobserved traffic flow can occur not only at the split but also through vehicles exiting in the studied direction.
@@
Previously, we evaluated our method on test data from the same section used for training in Subsection 6.2, demonstrating temporal transferability with spatial consistency. This scenario applies when sensors such as radars can be temporarily installed for training data collection and then removed to reduce long-term maintenance costs. Here, we investigate spatial transferability by training on one section (e.g., N1-IN) and testing on test days of different sections (e.g., N3-IN) without fine-tuning. This reflects scenarios where radar installation is infeasible or costly on adjacent sections but queue estimation remains necessary. Our adaptive design (Subsection 5.2) enables model transfer across sections of varying lengths.
Figure 10: Box plots of absolute estimation errors across all test days using different training sections.
@@
(b) Q-Net trained on N1-IN tested on N3-OUT..
Figure 11: Queue length estimation for section N3-IN and N3-OUT on 2023-11-14 using a model trained on N1-IN without fine-tuning.
Figure 10 presents absolute estimation errors under three training strategies: models trained individually on N1-IN, N3-IN, and N3-OUT, evaluated across all sections. As expected, each section achieves best performance when the model is trained on the same section, confirming effective capture of location-specific traffic dynamics. Performance on N3-IN and N3-OUT is similar regardless of which of these two sections is used for training, attributable to their spatial proximity and shared characteristics.
Interestingly, the model trained on N1-IN generalizes better to N3-IN and N3-OUT than vice versa. This asymmetry stems from N1-IN’s simpler traffic conditions and cleaner detector measurements, which enable learning of more accurate system dynamics representations with better convergence and stability. In contrast, N3-IN and N3-OUT exhibit higher complexity due to diverging over-ground branches that rejoin downstream (Figure 9). These branches create measurement challenges as downstream detector counts include vehicles not part of the initial section inflow, introducing unobserved traffic flows at both boundaries. The increased noise and structural ambiguities reduce the ability of data-driven models to learn reliable patterns, particularly when training data is limited.
Finally, Figure 11 visualizes the performance on N3-IN and N3-OUT for the test day 2023-11-14, using Q-Net trained on N1-IN. Our model successfully captures multiple queue peaks throughout the day, with the estimated queue (green line) closely following the ground truth (blue line) during most queue formation and dissipation phases. This demonstrates effective queue trend estimation for both unseen sections and time periods. Two noteworthy observations emerge from the results. First, the ground truth queue never drops to 0m, indicating potential radar malfunction since continuous queues even during early morning hours are unrealistic. Second, during afternoon peak periods when queues stabilize, Q-Net provides higher estimates than ground truth. This conservative bias likely reflects radar detection limitations and potential spillback to upstream intersections, but is acceptable since overestimation helps ensure activation of appropriate traffic control strategies.
@@ 6.4 Online estimation for real-time traffic control
Previously, in Subsection 4.1, we introduced the control input formulation based on queue length changes, which relies on estimating a fixed unobserved traffic flow rate via linear regression using cumulative net flow data over an entire day. While this approach is effective for ex post assessment and performance evaluation, it is not suitable for real-time applications where timely updates are essential. In this subsection, we extend our method to support real-time traffic control by dynamically updating the unobserved traffic flow rate during operation.
The underlying idea is straightforward: rather than computing a static unobserved traffic flow rate from full-day data, we continuously estimate it in real-time using cumulative flow data from the start of the day up to the current moment. Figure 12 compares the performance of the fixed (full-day) rate with the real-time adaptive version. Both methods are capable of capturing the general trend of queue dynamics. However, the real-time variant exhibits slightly reduced accuracy, which is expected given its reliance on shorter-term data that may be more susceptible to temporary fluctuations. Nonetheless, the ability to adapt online makes this approach more suitable for real-time traffic monitoring and control applications.
(a) Queue length estimation on 2023-11-14 (Tuesday).
(b) Queue length estimation on 2023-11-22 (Wednesday).
Figure 12: Comparison of queue length estimation for section N1-IN using two approaches for identifying unobserved traffic flow rate. The first estimates a fixed rate from full-day data via linear regression, while the second dynamically learns and adapts the rate in real time.
@@ 6.5 Ablation study
In this subsection, we examine the contributions of different components of our model through ablation analysis. Abewickrema et al. (2023) employed a Kalman filter for queue estimation, assuming that the maximum queue length remains constant over each signal cycle in the state evolution equation. This assumption motivates the design of our first variant, Q-Net w/o $u_{t}$, which excludes the control input derived from cumulative vehicle counts and instead assumes that the queue length remains unchanged over time. This variant relies solely on aFCD as input. We also introduce Q-EKF, in which the KalmanNet module is replaced by a standard Extended Kalman Filter (EKF), allowing us to isolate the benefit of using a learned observation update in the presence of nonlinear and partially known system dynamics.
@@
Default 71.65 49.51 48.60 5.29 36.94 79.43 26.79 68.04 17.91
Q-Net w/o $u_{t}$ 81.63 53.27 54.10 9.11 39.70 86.46 16.02 59.79 16.95
Q-EKF 229.79 189.27 292.12 175.09 180.39 446.28 117.27 121.00 49.83
@@
Figure 13: Queue length estimation results for different model variants on 2023-11-14 for section N1-IN.
The performance comparison of the model variants is presented in Table 4. The default Q-Net demonstrates the best overall accuracy and outperforms all variants during the morning peak period. However, its performance is slightly worse than Q-Net w/o $u_{t}$ during the afternoon peak hours. This is likely because afternoon peaks are typically characterized by consistently long queues with minimal variation, reducing the benefit of incorporating queue change dynamics. Additionally, Q-EKF yields significantly higher errors, underscoring the importance of a learnable Kalman gain. This result highlights the advantage of using a data-driven approach to dynamically track error covariances, as implemented in KalmanNet.
In Figure 13, we visualize the queue length estimation results on 2023-11-14 produced by the model variants and our default configuration. Q-Net w/o $u_{t}$ exhibits a noticeable delay in capturing the queue formation during the morning peak period. This lag is likely due to noise in the aFCD data, which becomes more impactful in the absence of control input. Q-EKF, while able to follow some general trends, fails to accurately capture both the queue formation and dissipation phases. These findings underscore the limitations of fixed-gain approaches like EKF in the context of highly dynamic urban traffic, and collectively support the design choices underlying our proposed Q-Net.

### v2

@@ 6.1 Performance
Table 3: Performance comparison of the proposed and baseline models when trained and tested on N1-IN. Improvement is over best baseline model.
@@
OSD 200.17 121.56 95.48 122.94 71.24 118.53 349.54 269.32 58.02
ISC 191.18 131.22 94.03 165.84 108.07 135.33 297.51 241.49 50.3
Tree-RPROP 114.25 67.96 65.11 111.17 58.5 91.74 101.37 75.49 19.43
Q-Net (ours) 86.77 58.14 56.22 78.62 47.62 74.64 88.36 69.25 17.41
Improvement 24.05% 14.44% 13.65% 29.28% 18.61% 18.64% 12.84% 8.26% 10.37%
Figure 8: Box plots of absolute estimation errors for baseline models (OSD, ISC) and Q-Net for section N1-IN. All p-values between the error distributions of baseline methods and our approach are less than $10^{-5}$, indicating statistically significant differences.
@@
(d) Q-Net vs Tree-RPROP.
Figure 9: The aFCD visualization and queue length estimation comparison among baselines (OSD, ISC, and Tree-RPROP), and our proposed method for section N1-IN on 2023-11-14. Results are generated using the model trained with random seed 42.
Quantitative analysis. Model performance comparison is presented in Table 3, with absolute error distributions shown in Figure 8. Q-Net consistently outperforms all baseline approaches, achieving significantly lower mean absolute errors and reduced error variance across all periods. Among the baselines, OSD and ISC perform worst, as their direct reliance on aFCD data makes them highly sensitive to aFCD errors. Despite Tree-RPROP employing more data sources, Q-Net consistently outperforms it. This is attributable to Q-Net’s novel state-space model, which is designed to accommodate unobserved traffic flow situations, and its AI-augmented Kalman filter, which effectively learns time-varying noises in the queue model.
A notable observation from Table 3 is the higher MAPE for all methods during the morning peak hours. By examining the ground-truth data, we saw that queues during this period do not consistently extend across the entire 07:00–09:00 window. As MAPE expresses the error relative w.r.t. the ground truth value, shorter queues produce larger percentage errors. As Figure 8 illustrates, all methods exhibit smaller errors during weekends and morning peaks, as queues are less pronounced during these periods, resulting in shorter queues that are easier to predict.
Qualitative analysis. A visual case study for 2023-11-14 is illustrated in Figure 9. Baseline OSD exhibits sharp, erratic fluctuations, while ISC suffers from discontinuities and generates multiple queue estimates for a single time step. Even the strongest baseline, Tree-RPROP, produces many unrealistic maximum queues, a result of its reliance on a fixed loop detector occupancy threshold in noisy conditions. In contrast, Q-Net successfully captures both the trend and variations, producing a smooth and stable estimation.
The poor performance of OSD and ISC is rooted in their sensitivity to aFCD errors, which are illustrated in Figure 9(a). These errors propagate in three ways: (a) erroneous high speeds recorded within the queue during congestion (e.g., 15:00–19:00), causing both baselines to place the queue end incorrectly in the middle of the section; (b) high speeds reported near the stop line (the first FCD segment), likely due to sparse probe vehicle coverage or artifacts from aggregation, which leads baselines to estimate a zero or near zero queue despite congestion; and (c) isolated, anomalous low-speed patches that cause ISC to generate multiple queue estimates for a single time step.
Q-Net’s architecture is specifically designed to overcome these challenges. Its AI-augmented Kalman filter dynamically tracks noise covariance, producing a robust estimate even under these noisy input conditions. Focusing on the period from 13:30 to 14:00, we observe a queue formation delay in the aFCD. Despite this, Q-Net accurately tracks the onset of the queue, thanks to the initial estimate derived from the loop detector cumulative count data, which reflects queue changes more promptly, thereby compensating for the aFCD lag.
@@ 6.2 Transferability
In Subsection 6.1, we demonstrated temporal transferability, where the model was trained and tested on the same section but using different time periods. This scenario applies when sensors such as radars can be temporarily installed for training data collection and then removed to reduce long-term maintenance costs. However, radar installation can be infeasible or costly on adjacent sections. This subsection investigates spatial transferability: the model is trained on one section and tested on test days of different sections without fine-tuning. Our adaptive design (Subsection 5.2) enables model transfer across sections of varying lengths.
@@
For this analysis, we introduce two additional sections from the Rotterdam urban area, both adjacent to N1-IN: N3-IN and N3-OUT, as shown in Figure 10. Both sections are located underground and feature an above-ground split after vehicles exit their upstream intersections. These diverging flows are eventually merged at the downstream intersection. The upstream intersection of N3-IN includes lanes with combined directions (through and right-turn), meaning unobserved traffic flow can occur not only at the split but also through vehicles exiting in the studied direction. Our experiment consists of training models individually on N1-IN, N3-IN, and N3-OUT and then evaluating each model across all three sections.
Figure 11: Box plots of absolute estimation errors across all test datasets using different training sections.
Figure 11 presents the performance for all training and testing combinations. As expected, each section achieves its best performance when the model is trained on the same section, confirming that the model effectively captures location-specific traffic dynamics. Performances on N3-IN and N3-OUT are close regardless of which of these two sections is used for training, attributable to their spatial proximity and shared infrastructure characteristics.
Interestingly, the model trained on N1-IN generalizes better to N3-IN and N3-OUT than vice versa. N1-IN features a simpler traffic infrastructure layout, the unobserved flow from its upstream split enables the model to learn robust noise representations that transfer effectively to N3-IN and N3-OUT. In contrast, N3-IN and N3-OUT exhibit additional complexity: diverging over-ground branches that rejoin downstream (Figure 10) introduce unobserved traffic at both boundaries, and downstream detector counts mix vehicles not part of the initial inflow. The increased noise and structural ambiguities make it difficult for the model to separate uncertainties from multiple sources, particularly when training data are limited. This leads to reduced transfer performance to N3-IN, where only upstream uncertainty is pronounced.
@@
(b) Q-Net trained on N1-IN tested on N3-OUT on 2023-11-14.
Figure 12: Queue length estimation for section N3-IN and N3-OUT using a model trained on N1-IN without fine-tuning (random seed 42).
Finally, Figure 12 visualizes the performance on N3-IN and N3-OUT for the test day 2023-11-14, using Q-Net trained on N1-IN. Our model successfully captures multiple queue peaks throughout the day, with the estimated queue (green line) closely following the ground truth (blue line) during most queue formation and dissipation phases. This demonstrates effective queue trend estimation for both unseen sections and time periods. Two noteworthy observations emerge. First, the ground truth queue never drops to 0m, indicating potential radar malfunction since continuous queues even during early morning hours are unrealistic. Second, during afternoon peak periods, when queues stabilize, Q-Net produces higher estimates than the ground truth. From a traffic management perspective, this conservative bias is acceptable, since overestimations have negligible operational impact once the minimum queue required to trigger traffic control is reached.
@@ 6.3 Online estimation for real-time traffic control
In Subsection 4.1, we introduced the control input formulation based on queue length changes, which relies on estimating a fixed unobserved traffic flow rate via linear regression using cumulative net flow data over an entire day. While this approach is effective for ex post assessment and performance evaluation, it is not suitable for real-time applications where timely updates are essential. In this subsection, we evaluate the ability of our method in supporting real-time traffic control by dynamically updating the unobserved traffic flow rate during operation.
The underlying idea is straightforward: rather than computing a static unobserved traffic flow rate in Equation (16) from full-day data, we estimate it in real-time using cumulative flow data from the start of the day up to the current moment. The same hyperparameters are used as in the default setting.
Figure 13 compares the performance of the fixed (full-day) rate with the online adaptive version. Both methods capture the general trend of queue dynamics. The online variant shows slightly lower accuracy—for instance, it exhibits a greater queue formation delay between 11:30–12:30 and errors in queue dissipation around 10:30–11:00 on 2023-11-22. This is expected, as its reliance on shorter-term data that are more susceptible to temporary fluctuations. Nevertheless, Table 4 shows that the average performance of the online version and the default Q-Net are close, and the online model’s adaptability makes it more suitable for real-time traffic monitoring and control applications.
Table 4: Performance comparison of Q-Net and its onlie version when trained and tested on N1-IN.
All day Morning peak Afternoon peak
RMSE MAE MAPE RMSE MAE MAPE RMSE MAE MAPE
Default Q-Net 86.77 58.14 56.22 78.62 47.62 74.64 88.36 69.25 17.41
Online Q-Net 85.33 56.74 56.34 76.24 46.93 74.88 81.07 62.6 15.56
Figure 13: Comparison of queue length estimation for section N1-IN on 2023-11-22 (Wednesday) using default and online Q-Net. The first estimates a fixed rate from full-day data via linear regression, while the second dynamically learns and adapts the rate in real time.
@@ 6.4 Ablation study
Default 86.77 58.14 56.22 78.62 47.62 74.64 88.36 69.25 17.41
90.2 59.75 56.81 83.31 49.24 73.24 87.54 68.67 17.27
Q-Net w/o $u_{t}$ (-3.9%) (-2.8%) (-1.1%) (-6.0%) (-3.4%) (+1.9%) (+0.9%) (+0.8%) (+0.8%)
217.58 175.78 259.58 213.45 165.46 353.47 135.55 111.59 34.39
Q-EKF (-150.7%) (-202.3%) (-361.7%) (-171.5%) (-247.5%) (-373.5%) (-53.4%) (-61.1%) (-97.5%)
90.12 60.26 58.0 88.1 51.4 83.91 89.78 71.3 16.63
Q-Net w/o group (-3.9%) (-3.6%) (-3.2%) (-12.1%) (-8.0%) (-12.4%) (-1.6%) (-3.0%) (+4.5%)
As our key innovations lie in the state-space model formulation, the data-driven Kalman filtering approach, and the grouped measurement strategy, this subsection presents an ablation analysis to examine their respective contributions. For the state-space model, inspired by Abewickrema et al. (2023), which assumes that the maximum queue length remains constant over each signal cycle in the state evolution equation, we design the first variant, Q-Net w/o $u_{t}$. This model assumes that the queue length remains unchanged over time by excluding the control input $u_{t}$ in Equation (11): $x_{t}=x_{t-1}+w_{t}$. Since $u_{t}$ is derived from cumulative vehicle counts, removing this term makes the variant rely solely on aFCD data. The second variant, Q-EKF, replaces the KalmanNet module with a standard Extended Kalman Filter (EKF). This allows us to isolate the benefit of using a learned state update in the presence of nonlinear and partially known system dynamics. The third variant, Q-Net w/o grouping, removes our proposed grouping strategy (Equation (33)). Consequently, this variant is trained as location-specific without spatial transferability. The comparison is presented in Table 5 and Figure 14).
@@
Figure 14: Queue length estimation results for different model variants on 2023-11-22 for section N1-IN.
State-space formulation and control input ($u_{t}$). The inherent value of the state-space model is demonstrated by comparing Q-Net w/o $u_{t}$ with the baselines from Subsection 6.1 (Table 3). Despite using the same aFCD-only data, Q-Net w/o $u_{t}$ achieves substantial improvement, highlighting the value of state-space modeling for capturing the temporal evolution of queues.
The contribution of the control input ($u_{t}$) is shown by comparing Q-Net w/o $u_{t}$ to the default Q-Net. The Q-Net w/o $u_{t}$ variant, relying only on aFCD, is more susceptible to signal lag and exhibits noticeable delays in capturing queue formation (e.g., 11:30-13:00, Figure 14). The default Q-Net, by incorporating the flow-count-based $u_{t}$, effectively mitigates this lag and better tracks these dynamic queueing phases.
Interestingly, Q-Net w/o $u_{t}$ performs slightly better in Table 5 during afternoon peaks, when queues are long and stable. This means that the true queue length change ($u_{t}$) is minimal and often near zero. The variant’s assumption of $u_{t}\approx 0$ is thus a valid approximation of this specific physical state. In this condition, the full model’s $u_{t}$ term, estimated from noisy cumulative counts, adds little corrective information and leads to minor performance degradation. This nuance confirms the overall value of the control input, which is more critical for tracking queue formation and dissipation.
Data-driven Kalman filtering. The contribution of our data-driven filter is straightforward. The Q-EKF variant is the worst-performing model across all metrics, as it fails to accurately track queue formation and dissipation (Figure 14). This poor performance highlights the critical importance of a learnable Kalman gain. The EKF’s reliance on static, predefined noise covariances is a fundamental limitation, and this result confirms it is insufficient for tracking the complex noise in highly dynamic real-world urban conditions.
Grouped measurement strategy. The Q-Net w/o group variant, trained as a location-specific model, achieves performance worse than the default Q-Net (Table 5). This implies that our default model pays almost no accuracy penalty for the structural constraints that enable spatial transferability. The default model’s improvement suggests that focusing on neighboring segments provides valuable contextual information. This result justifies our grouping strategy, as it achieves its primary goal (spatial transferability, as shown in Subsection 6.1) without compromising single-location performance.
7 Discussion
We proposed Q-Net, a computationally and data-efficient framework for queue length estimation. Our problem setting reflects a general and practically relevant scenario in which traffic flow is only partially observed. While loop detectors can be flexibly positioned along the roadway, our method remains applicable by accounting for the time lag between vehicle detection and arrival at the stop line, though performance may vary depending on detector placement and traffic conditions. Q-Net relies solely on two widely available and privacy-preserving data sources: cumulative vehicle counts from loop detectors and aggregated floating car data (aFCD). Together, these aspects highlight the broad applicability of the proposed Q-Net.
Note on architectural design: Q-Net is built on a state-space formulation tailored for queue length estimation, with a neural network–enhanced filtering algorithm that enables adaptive learning. Loop detector data are used to derive queue change (the control input), which describes the temporal evolution of queues. As detailed in Subsection 4.1, this control input is derived through affine transformation and Fourier-based band-pass filtering. To improve the estimation of queue change, one could design specialized neural architectures for this component. We experimented with long short-term memory (LSTM) and attention mechanisms; however, empirical results showed that our current formulation yielded superior performance. Under limited data conditions, the model struggles to learn queue change dynamics and noise uncertainties at the same time. Although it is possible to focus solely on queue change estimation and subsequently infer queue length, such an approach lies beyond the scope of this study.
In the filtering process, we apply a projection operation to ensure that the estimated queue length remains within physically plausible bounds. While this sacrifices some mathematical optimality, such ad hoc constraints are widely adopted in engineering practice (Simon, 2010; Prakash et al., 2014) and demonstrated good performance in our experiments.
Regarding the AI-augmented filtering process, as described in Subsection 5.1, FC layers are typically followed by ReLU activations to produce non-negative feature representations. However, this architecture does not provide theoretical guarantees of positive semi-definiteness for the estimated covariance matrices or strict non-negativity for scalar variance estimates. In particular, outputs from the GRU layers are unconstrained, and the Kalman gain is computed via a two-layer network without explicit non-negativity enforcement on the final output. While techniques such as Cholesky decomposition or eigenvalue regularization can impose these constraints, we observed that unconstrained training achieves sufficient accuracy for our application. This design choice offers the network flexibility to learn effective representations for the filtering task while avoiding the computational overhead and potential optimization challenges associated with hard constraints on matrix properties.
Evaluation: We recognize two primary limitations in our dataset. First, while the data is real-world, it is limited to 16 days within a single month. This temporal constraint is, however, common in the queue estimation domain (Liu et al., 2009; Abewickrema et al., 2023; Luo et al., 2022). While one could use simulation tools such as SUMO (Sengupta et al., 2021b) or AIM-SUN (Abewickrema et al., 2025) to generate more data, we did not adopt this approach. Modeling the unobserved traffic flows in a high-fidelity manner is a complex research problem in itself and falls beyond the scope of this research. Second, our spatial transferability analysis was limited to N3-IN and N3-OUT. These sections are adjacent to N1-IN and therefore not fully independent. Although their geometric and flow differences introduce some variation, generalization to entirely separate networks remains an open question.
Expanding this analysis to more diverse and independent sites is an important next step. Future work could also explore alternative training paradigms, such as curriculum learning, and validate the approach across a wider range of urban environments to strengthen generalizability. Furthermore, enhancing Q-Net with graph-based learning could capture spatial correlations across intersections, enabling more expressive evolution models for queue dynamics. Integrating additional contextual features, such as traffic signal timings or weather conditions, may further boost accuracy. Ultimately, combining Q-Net with traffic control and optimization strategies opens new avenues for intelligent, data-driven traffic management systems.